Dearbrook University offers 16 sections of contemporary history. If Dr. Jains adds 7 of those sections to his schedule, what percentage of the contemporary history sections did Dr. Jains add to his schedule? Round the answer to the nearest tenth.
1
Expert's answer
2019-09-13T10:27:14-0400

No of the contemporary history sections = 16

No of sections added to Jains' schedule = 7

percentage of the sections added to the schedule = (No of sections added to the schedule) / No of the sections)*100

= 7/16 * 100 = 0.4375* 100 = 43.75 %

Rounding to the nearest renth = 43.8 percent

The percentage of the contemporary history sections added by Dr. Jains to his schedule was 43.8